Two coins are tossed simultaneously $500$ times,and we get:
Two heads : $105$ times
One head : $275$ times
No head : $120$ times
Find the probability of occurrence of each of these events.

Vedclass pdf generator app on play store
Vedclass iOS app on app store
(N/A) Let us denote the events of getting two heads,one head,and no head by $E_1, E_2,$ and $E_3$ respectively.
The total number of trials is $500$.
For $E_1$ (two heads): $P(E_1) = \frac{105}{500} = 0.21$
For $E_2$ (one head): $P(E_2) = \frac{275}{500} = 0.55$
For $E_3$ (no head): $P(E_3) = \frac{120}{500} = 0.24$
Observe that $P(E_1) + P(E_2) + P(E_3) = 0.21 + 0.55 + 0.24 = 1.0$. These events cover all possible outcomes of the trial.

Explore More

Similar Questions

An insurance company selected $2000$ drivers at random in a particular city to find a relationship between age and accidents. The data obtained are given in the following table:
Age of drivers (in years) $0$ accidents $1$ accident $2$ accidents $3$ accidents Over $3$ accidents
$18-29$ $440$ $160$ $110$ $61$ $35$
$30-50$ $505$ $125$ $60$ $22$ $18$
Above $50$ $360$ $45$ $35$ $15$ $9$

Find the probabilities of the following events for a driver chosen at random from the city:
$(i)$ Being $18-29$ years of age and having exactly $3$ accidents in one year.
$(ii)$ Being $30-50$ years of age and having one or more accidents in a year.
$(iii)$ Having no accidents in one year.

The record of a weather station shows that out of the past $250$ consecutive days,its weather forecasts were correct $175$ times.
$(i)$ What is the probability that on a given day it was correct?
$(ii)$ What is the probability that it was not correct on a given day?

Fifty seeds were selected at random from each of $5$ bags of seeds,and were kept under standardised conditions favourable to germination. After $20$ days,the number of seeds which had germinated in each collection were counted and recorded as follows:
Bag $1$ $2$ $3$ $4$ $5$
Number of seeds germinated $40$ $48$ $42$ $39$ $41$

What is the probability of germination of
$(i)$ more than $40$ seeds in a bag?
$(ii)$ $49$ seeds in a bag?
$(iii)$ more than $35$ seeds in a bag?

Ask all the students in your class to write a $3-$digit number. Choose any student from the room at random. What is the probability that the number written by her/him is divisible by $3$? Remember that a number is divisible by $3$,if the sum of its digits is divisible by $3$.

In a cricket match,a batswoman hits a boundary $6$ times out of $30$ balls she plays. Find the probability that she did not hit a boundary.

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo